OpenStudy (anonymous):

Find the product: -4x (-7x^3 + 4 x 2 + 8x – 6)

OpenStudy (anonymous):

-4x (-7x^3 + 4 x 2 + 8x – 6) Okay, you take the -4x and multiply it by everything in there, so we'll start with the -7x^3: -4x * -7x^3 = 28x^4 (negative times a negative is a positive and when multiplying like exponents, you add the exponents so x * x^3 = x^4 Now, how do you do -4x * 4?

OpenStudy (anonymous):

-16x

OpenStudy (anonymous):

but this is college and adding and subtractin polynomials

OpenStudy (anonymous):

@melbel is on the right track. following what they said,  Can u multiply that out?